# Activated LoRA: Fine-tuned LLMs for Intrinsics

### v2

@@ Abstract
Low-Rank Adaptation (LoRA) has emerged as a highly efficient framework for finetuning the weights of large foundation models, and has become the go-to method for data-driven customization of LLMs. Despite the promise of highly customized behaviors and capabilities, switching between relevant LoRAs in a multiturn setting is highly inefficient, as the key-value (KV) cache of the entire turn history must be recomputed with the LoRA weights before generation can begin. To address this problem, we propose Activated LoRA (aLoRA), which modifies the LoRA framework to only adapt weights for the tokens in the sequence after the aLoRA is invoked. This change crucially allows aLoRA to accept the base model’s KV cache of the input string, meaning that aLoRA can be instantly activated whenever needed in a chain without recomputing the cache. This enables building what we call intrinsics, i.e. highly specialized models invoked to perform well-defined operations on portions of an input chain or conversation that otherwise uses the base model by default. We train a set of aLoRA-based intrinsics models, demonstrating competitive accuracy with standard LoRA while achieving significant inference benefits. The codebase is at https://github.com/IBM/activated-lora.
@@ 1 Introduction
The rapid adoption of large language models (LLMs) has catalyzed significant advancements in natural language processing tasks, from text generation to knowledge extraction. However, adapting these models to specific tasks or domains often demands fine-tuning their immense parameter space, a process that is computationally expensive and difficult to scale. Low-Rank Adaptation (LoRA) (Hu et al., 2021) has addressed these challenges by introducing a parameter-efficient method for fine-tuning (Houlsby et al., 2019), enabling highly customized model behavior without the need to retrain or modify the entire model. By optimizing a small subset of low-rank matrices, LoRA has emerged as a lightweight and effective alternative for task-specific customization, particularly for large foundation models. Yet, while LoRA excels in static or single-task scenarios, it presents inefficiencies when applied in multiturn interactions, where dynamic switching between multiple LoRA configurations is required in agentic or reasoning pipelines.
@@
This inefficiency arises from the necessity to recompute the representation of all tokens prior to generation when switching between LoRA weights. In the case of the popular attention mechanism introduced by Vaswani et al. (2017), such a representation takes the form of the key-value (KV) cache of such prior tokens, but more generally, it is a problem that reoccurs even in different architectures; for example, in a state-space model, the state representation of the tokens prior to generation would need to be recomputed if the matrices $A$ and $B$ are updated by a low rank correction. This recalculation introduces significant latency, GPU memory, and computational overhead that all scale with the length of the context that must be prefilled. This limits LoRA’s usability in scenarios where rapid transitions between specialized behaviors are essential.
If only a few specialized behaviors are needed, all needed tasks can in principle be trained into the LoRA while retaining as much of the base model behavior as possible (here the LoRA would be used for all parts of the pipeline), but this solution has major limitations. In particular, by modifying the weights of the model for model calls that are best handled by the base model, performance degradation is likely on those tasks. Additionally, multi-task LoRA training is significantly more difficult than training a LoRA for a single task, and this approach is inherently non-modular, requiring complete retraining to incorporate new intrinsics abilities.
Addressing this inability to reuse the base model representation of the tokens prior to generation, we introduce Activated LoRA (aLoRA), a novel extension of the LoRA framework designed to dynamically adapt the model’s weights for tokens encountered after activation. By decoupling the adaptation process from the need to recompute the input’s representation, aLoRA facilitates instantaneous switching between specialized models—or “intrinsics”—while maintaining seamless interaction with the base model. This innovation not only reduces computational costs but also unlocks new possibilities for deploying highly modular, task-specific behaviors within complex workflows.
1.1 Background
We will illustrate the concept behind activated LoRA in the context of the ubiquitous attention mechanism introduced by Vaswani et al. (2017).
The Attention Mechanism
Recall that the attention mechanism in each attention layer in LLMs (see Figure 1 for a typical architecture) generally takes the form
where $d_{k}$ is the dimension and $Q,K,V$ are concatenated queries, keys, and values for the input tokens:
$Q=XW^{Q},\quad K=XW^{K},\quad V=XW^{V}$ (2)
where $W^{Q}$, $W^{K}$, and $W^{V}$ are weight matrices.
Figure 1: A generic LLM architecture.
LLM inference and the KV cache
LLMs generate tokens autoregressively one at a time. While we do not write out the full equations here, note that generating this next token only requires computing the last row of this attention matrix if the keys and values of all prior tokens are precomputed (e.g. from when they were generated themselves or prefilled). This is because the last row of the attention matrix only uses the last row of $Q$, but the entire $K$ and $V$ matrices. These precomputed $K$ and $V$ values are called the KV cache. This observation creates significant speedups for LLM inference and is the key to all modern LLM inference engines.
LoRA
LoRA adapts the attention weights $W^{Q}$, $W^{K}$, and $W^{V}$ by replacing them with $W^{Q}+\Delta^{Q}$, $W^{K}+\Delta^{K}$, and $W^{V}+\Delta^{V}$, where $\Delta^{Q},\Delta^{K},\Delta_{V}$ are rank $r$ matrices. This yields
This lowers the number of parameters, making finetuning significantly more efficient (Hu et al., 2021). If the LoRA is active for the entire chain, then the KV cache-based inference applies and generation is efficient. If, on the other hand, any part of the input was generated or prefilled by the base model or another LoRA, then the KV cache for the input must be recomputed.
2 Intrinsics and late prompts
@@
Within the discipline of software engineering, intrinsics are generally useful functions that are built into a programming language whose implementation can be optimized by a compiler. We define an LLM intrinsic to be a capability that can be invoked through a well defined API that is reasonably stable and independent across model generations and families of how the intrinsic is implemented. Performance metrics such as accuracy, latency and throughput may vary significantly across such implementations.
@@
The concept of activated LoRA takes an opinionated view on how such differentiated performance could be attained. As inspiration, we note that instructions in LLMs can appear in many places in a prompt; Figure 2 illustrates an example of two such places: an “early prompting” case where the instruction precedes the content, or a “late prompting” case where the content precedes the instruction. In the latter, the instruction does not have to be revealed ahead of time, making the representation of the content potentially useful for many tasks. We note that these paradigms can take on many shapes; for example, early prompting could just as well be substituted by prefix tuning methods such as Li and Liang (2021) with the same overall outcome.
The reader may appreciate that a LoRA adapter, like the prefix tuning example above, shares some similarities with the early prompting concept; the LLM’s internal representation of the content is fit-for-purpose, specific to the LoRA adapter and not easily reusable. Late prompts, on the other hand, suffer from a potential disadvantage compared to early prompts since there is no opportunity for the content to be represented with the specific task in mind. Can we find a LoRA-like techology, that gives us the benefits of late prompts with none of the potential accuracy disadvantages?
In the following section, we will address this shortcoming, presenting our Activated LoRA method, which activates adapted weights only on tokens corresponding to the intrinsic instruction and generation.
@@ 3 Activated LoRA
Our Activated LoRA (aLoRA) framework addresses this shortcoming. Just as in LoRA, aLoRA adapts the attention weights $W^{Q}$, $W^{K}$, and $W^{V}$ by replacing them with $W^{Q}+\Delta^{Q}$, $W^{K}+\Delta^{K}$, and $W^{V}+\Delta^{V}$, where $\Delta^{Q},\Delta^{K},\Delta^{V}$ are rank $r$ matrices. The difference lies in how these adapted weights are used. We assume that the the default generation model for the chat is the base model, and that intrinsics only operate on these base model generations. As a result, we can assume that the base model has precomputed a KV cache for the input context (the blue region in Figure 2).
@@
The aLoRA architecture is designed to directly reuse this base model KV cache. In the attention mechanism (1), aLoRA only adapts the $Q,K,V$ matrices for tokens occurring after the start of the invocation sequence. Specifically, instead of (3) we have
$Q=\left[\begin{array}[]{c}X_{before}W^{Q}\\ X_{after}(W^{Q}+\Delta^{Q})\end{array}\right],\quad K=\left[\begin{array}[]{c}X_{before}W^{K}\\ X_{after}(W^{K}+\Delta^{K})\end{array}\right],\quad V=\left[\begin{array}[]{c}X_{before}W^{V}\\ X_{after}(W^{V}+\Delta^{V})\end{array}\right],$ (4)
@@
where $X_{before}$ and $X_{after}$ are the portions of $X$ coming before and after the aLoRA model is invoked. Note that if $X_{before}$ is associated with tokens generated or prefilled by the base model, then $X_{before}W^{K}$ and $X_{before}W^{V}$ are already in the KV cache and do not need to be recomputed. Similarly, any tokens that have been prefilled or generated by the aLoRA model have keys and values processed by the adapted weights, so $X_{after}(W^{K}+\Delta^{K})$ and $X_{after}(W^{V}+\Delta^{V})$ are also available (except for the current token being generated). As a result, the aLoRA architecture can seemlessly reuse the existing base model KV cache as well as continue to maintain its own KV cache as it generates.
Note that the late prompting framework for intrinsics is highly amenable to supervised finetuning of aLoRA via data collators, and furthermore ensures that the tokens affected by the aLoRA are isolated to only the answer. As a result, once the aLoRA is turned off and the base model resumes generation, the base model can either resume from its own cache where the aLoRA began, or prefill the (usually short) answer from the aLoRA if needed.
3.1 Invocation
@@
We found it useful to demarcate the activation of the aLoRA adapter via an invocation token sequence. While the aLoRA will often be invoked programmatically by the runtime, this design in principle allows for the base model (or other intrinsics) to call the aLoRA model themselves. In our current implementation, except for the 1-token experiments below, the aLoRA weights are activated at the start of the invocation sequence, since it is typically inserted by the runtime and may not have a KV cache precomputed. Note that the invocation sequence is optional in principle, i.e. it can be an empty string.
The invocation sequence has several additional benefits. The invocation sequence can be designed to
  • Conform to the chat template (for instance by making the aLoRA response its own turn with a specialized role)
  • Provide a short prompt to aid the learning process (the name of the intrinsic may be enough)
  • Give the adapter weights more tokens to process the input prior to generating the output, possibly improving performance.
3.2 Training and Huggingface Implementation

### v4

@@ Abstract
Low-Rank Adaptation (LoRA) has emerged as a highly efficient framework for finetuning the weights of large foundation models, and has become the go-to method for data-driven customization of LLMs. Despite the promise of highly customized behaviors and capabilities, switching between relevant LoRAs in a multiturn setting is inefficient, as the key-value (KV) cache of the entire turn history must be recomputed with the LoRA weights before generation can begin. To address this problem, we propose Activated LoRA (aLoRA), an adapter architecture which modifies the LoRA framework to only adapt weights for the tokens in the sequence after the aLoRA is invoked. This change crucially allows aLoRA to accept the base model’s KV cache of the input string, meaning that aLoRA can be instantly activated whenever needed in a chain without recomputing the cache. This enables building what we call intrinsics, i.e. specialized models invoked to perform well-defined operations on portions of an input chain or conversation that otherwise uses the base model by default. We train a set of aLoRA-based intrinsics models, demonstrating competitive accuracy with standard LoRA while achieving significant inference benefits. The codebase is at https://github.com/IBM/activated-lora.111pip install alora
@@ 1 Introduction
The rapid adoption of large language models (LLMs) has catalyzed significant advancements in natural language processing tasks, from text generation to knowledge extraction. However, adapting these models to specific tasks or domains often demands fine-tuning their immense parameter space, a process that is computationally expensive and difficult to scale. Low-Rank Adaptation (LoRA)(Hu et al., 2021) has addressed these challenges by introducing a parameter-efficient method for fine-tuning (Houlsby et al., 2019), enabling highly customized model behavior without the need to retrain or modify the entire model. By optimizing a small subset of low-rank matrices, LoRA has emerged as a widely-used lightweight and effective alternative for task-specific customization, particularly for large foundation models such as LLMs, and popular LoRA finetuning services are offered by both corporate and open-source LLM providers (OpenAI, 2024; TogetherAI, 2024; Predibase, 2024). While large models perhaps have less need of finetuning, small models continue to see strong benefits from LoRA adapters Wang et al. (2025); Danilevsky et al. (2025).
Motivated by this, significant work has been done to enable highly efficient serving of and inference with LoRAs, e.g. Sheng et al. (2023); Brüel-Gabrielsson et al. (2024), with vLLM Kwon et al. (2023) incorporating and further optimizing many of these algorithms in its popular, state-of-the-art inference engine. Yet, while LoRA excels in static or single-task scenarios, in modern applications a wide mix of skills are typically needed, e.g. in multiturn interactions and agentic applications Wang et al. (2024); Yang et al. (2024). Crucially, LoRAs often must be carefully tuned to avoid forgetting/degradation of performance on the wide variety of tasks the base model already performs well on. In applications where dynamic switching between multiple specialized skills would be advantageous (e.g. agentic or reasoning pipelines), this strategy breaks down. Additionally, multi-task LoRA training is significantly more difficult than training a LoRA for a single task, and this approach would be inherently non-modular, requiring retraining to incorporate further abilities later on.
@@
In an ideal world, we would like to be able to seamlessly switch—within the same chat/agentic pipeline—between arbitrary sets of pretrained LoRA configurations for specialized tasks, while keeping the base model as-is for most interactions in the sequence. Unfortunately, the LoRA architecture is not well-suited to this regime, creating major inference inefficiencies. Specifically, for each switch between LoRA adapters, it becomes necessary to recompute the representation of all context tokens prior to generation with the LoRA. In the case of the popular attention mechanism introduced by Vaswani et al. (2017), such a representation takes the form of the key-value (KV) cache of such prior tokens, but more generally, it is a problem that reoccurs even in different architectures; for example, in a state-space model, the state representation of the tokens prior to generation would need to be recomputed if the matrices $A$ and $B$ are updated by a low rank correction. This recalculation introduces significant latency, GPU memory, and computational overhead that all scale with the length of the context that must be prefilled, especially as LLMs increasingly work with documents, files, or reasoning/agentic chains in the many (sometimes hundreds of) thousands of tokens Liu et al. (2025; 2024; 2023) (even millions for software engineering Jiang et al. (2025)). This limits LoRA’s usability in scenarios where rapid transitions between specialized behaviors are essential.
@@
In this work, we address this shortcoming, presenting our Activated LoRA method, which activates adapted weights only on tokens corresponding to a short intrinsic instruction and subsequent generation, leaving the weights for other context tokens unchanged. We elaborate on the setting we envision here. Within the discipline of software engineering, intrinsics are generally useful functions that are built into a programming language whose implementation can be optimized by a compiler. We define an LLM intrinsic to be a capability that can be invoked through a well defined API that is reasonably stable and independent across model generations and families of how the intrinsic is implemented. Performance metrics such as accuracy, latency, and throughput may vary significantly across such implementations.
@@
The concept of activated LoRA takes an opinionated view on how such differentiated performance could be attained. As inspiration, we note that instructions in LLMs can appear in many places in a prompt; Figure 1 illustrates an example of two such places: an “early prompting” case where the instruction precedes the content, or a “late prompting” case where the content precedes the instruction. In the latter, the instruction does not have to be revealed ahead of time, making the representation of the content (KV cache) potentially reusable for or from other tasks. We note that these paradigms can take on many shapes; for example, prefix tuning methods such as Li and Liang (2021) explicitly use an early prompting framework, tuning this early prompt. Recently, a set of real-world tasks (and trained adapters) conforming to the late-prompting intrinsics regime was presented in Danilevsky et al. (2025), in the context of building performant and robust RAG Lewis et al. (2020) pipelines. We believe that a vast array of intrinsics-style finetuning tasks can be created with applications throughout the agentic, chat, and reasoning spaces.
Note that a LoRA adapter, like the prefix tuning example above, shares the downsides of the early prompting concept; the LoRA-adapted LLM’s internal representation of the content (KV cache) is fit-for-purpose, specific to the LoRA adapter and not reusable by or from the base model or other adapters.
@@
Addressing this inability to reuse the base model representation of the tokens prior to generation, we introduce Activated LoRA (aLoRA), a novel extension of the LoRA framework designed to only adapt the model’s weights for tokens encountered after activation, allowing base model KV cache for prior context to be reused (Figure 2). By decoupling the adaptation process from the need to recompute the input’s representation, aLoRA facilitates instantaneous switching between specialized models (intrinsics) while maintaining seamless interaction with the base model. This innovation not only reduces computational costs, but also unlocks new possibilities for deploying highly modular, task-specific behaviors within complex workflows.
In our experiments, we demonstrate significant speedups for aLoRA vs LoRA on the state-of-the-art inference engine vLLM Zhu et al. (2023), and show that aLoRA adapters do not lose accuracy versus LoRA on a collection of benchmark finetuning tasks and a collection of real-world “intrinsics” tasks from Danilevsky et al. (2025).
(a)promptanswer eval1eval2prefill (base)gen (base)prefill (LoRA1)prefill (LoRA1) gen (LoRA1)prefill (LoRA2)prefill (LoRA2) gen (LoRA2)(b)promptanswer eval1eval2prefill (base)gen (base)gen (aLoRA1)gen (aLoRA2)
Figure 2: Computation and memory pattern of (a) LoRA vs. (b) aLoRA used as evaluators of an answer given by a base model. (1) prompt is input to the base model, which generates answer, (2) prompt + answer is input to the both intrinsics in parallel, which generate eval_1 and eval_2 respectively. Narrow rectangles denote tokens and wide rectangles denote the KV cache.
1.1 Background
Modern decoder-only causal LLMs are transformers containing a sequence of decoder layers, each of which typically contain MLP layers that operate on each token’s representation independently, and a causal attention mechanism that allows tokens to attend to representations of prior tokens in the same layer (see Figure 9 in the appendix). We review the ubiquitous softmax attention mechanism introduced by Vaswani et al. (2017) and its connection to the KV cache, as well as LoRA adapters. [^2] Extension of this to other forms of attention is straightforward, but we leave this treatment to future work. Further, note that LoRA and aLoRA adapters can be applied to MLP blocks as well, but since it is typical practice to not adapt the MLP blocks we focus the presentation on the attention blocks only.
Attention: Recall that the softmax attention mechanism in each attention layer takes the form
where $d_{k}$ is the dimension and $Q,K,V$ are concatenated queries, keys, and values for the tokens:
$Q=XW^{Q},\quad K=XW^{K},\quad V=XW^{V}$ (2)
where $W^{Q}$, $W^{K}$, and $W^{V}$ are weight matrices.
LLM inference and the KV cache: LLMs, using a causal attention mask Radford et al. (2019), generate tokens autoregressively one at a time. Suppose we have already generated tokens $11,\dots,t-$ 1 with corresponding hidden-state embeddings $x_{1},\dots,x_{t-1}$. Whenever these tokens were processed (either during prefill or generation) their keys and values were already computed:
$K_{1:t-1}$ $=[\,K_{1};\dots;K_{t-1}\,],V_{1:t-1}=[\,V_{1};\dots;V_{t-1}\,],\quad K_{i}=x_{i}W^{K},V_{i}=x_{i}W^{V}.$
Due to causal attention, tokens can only attend to prior tokens, and hence are not affected by subsequent generations and can be stored for use by future tokens. At generation step $t$, we therefore only need to compute the new query, key, and value for the latest token
$Q_{t}=x_{t}W^{Q},\quad K_{t}=x_{t}W^{K},\quad V_{t}=x_{t}W^{V},$ (3)
and then form the full cached matrices by appending $K_{1:t}=\bigl{[}K_{1:t-1};\,K_{t}\bigr{]},V_{1:t}=\bigl{[}V_{1:t-1};\,V_{t}\bigr{]}.$ Finally, the output hidden state for token $t$ is ($d_{k}$ is hidden dimension)
$h_{t}=\mathrm{Attention}\bigl{(}Q_{t},\;K_{1:t},\;V_{1:t}\bigr{)}=\mathrm{softmax}\!\Bigl{(}\tfrac{Q_{t}\,K_{1:t}^{\top}}{\sqrt{d_{k}}}\Bigr{)}\;V_{1:t}.$ (4)
Because the entire matrix of past keys $K_{1:t-1}$ and values $V_{1:t-1}$ is precomputed and stored in the “KV cache,” at each new step $t$ we only pay the cost of computing one new row for each of $Q$, $K$, and $V$, and performing a single-row softmax and weighted sum over the remaining cached $K$ and $V$. As shown in Vaswani et al. (2017); Radford et al. (2019), this reduces what would have been an $\mathcal{O}(t^{2}d_{k})$ full-matrix multiply down to only $\mathcal{O}(t\,d_{k})$ operations at each step, yielding massive speedups.
LoRA: LoRA adapts the attention weights $W^{Q}$, $W^{K}$, and $W^{V}$ by replacing them with $W^{Q}+\Delta^{Q}$, $W^{K}+\Delta^{K}$, and $W^{V}+\Delta^{V}$, where $\Delta^{Q},\Delta^{K},\Delta_{V}$ are rank $r$ matrices. This yields
This lowers the number of parameters, making finetuning significantly more efficient (Hu et al., 2021). If the LoRA is active for the entire chain, then the KV cache-based inference applies and generation is efficient. If, on the other hand, any part of the input was generated or prefilled by the base model or another LoRA, then $K$ and $V$ for the LoRA are different than the corresponding $K$ and $V$ for the base model, hence any existing base model KV cache cannot be used by the LoRA, meaning the KV cache for the entire (potentially very long) input must be recomputed. To clarify, this involves passing the input through all layers of the transformer in sequence–not just the adapted attention layers–since modifications to attention blocks modify the inputs to all downstream layers and tokens.
@@ 2 Activated LoRA
Just as in LoRA, our aLoRA architecture adapts the attention weights $W^{Q}$, $W^{K}$, and $W^{V}$ by replacing them with $W^{Q}+\Delta^{Q}$, $W^{K}+\Delta^{K}$, and $W^{V}+\Delta^{V}$, where $\Delta^{Q},\Delta^{K},\Delta^{V}$ are rank $r$ matrices. The difference lies in how these adapted weights are used. We assume that the the default generation model for the chat is the base model, and that intrinsics only operate on these base model generations. As a result, we can assume that the base model has precomputed a KV cache for the input context (the blue region in Figure 1).
@@
The aLoRA architecture is designed to match the base model keys and values on context tokens, allowing the adapter to reuse base model KV cache for those tokens, or, vice versa, the base model to reuse KV cache that the aLoRA adapter creates for its context inputs. Specifically, in the attention mechanism (1), aLoRA only adapts the $Q,K,V$ matrices for tokens occurring after the start of the invocation sequence. Let the token index for adapter activation be $t_{\mathrm{invoke}}$, and let the current token have index $t\geq t_{\mathrm{invoke}}$. Instead of (5) we then have
$Q\!=\!\!\left[\!\!\begin{array}[]{c}X_{1:t_{\mathrm{invoke}}-1}W^{Q}\\ X_{t_{\mathrm{invoke}}:t}(W^{Q}\!+\!\Delta^{Q})\end{array}\!\!\right],K\!=\!\!\left[\!\!\begin{array}[]{c}X_{1:t_{\mathrm{invoke}}-1}W^{K}\\ X_{t_{\mathrm{invoke}}:t}(W^{K}\!+\!\Delta^{K})\end{array}\!\!\right],V\!\!=\!\left[\!\!\begin{array}[]{c}X_{1:t_{\mathrm{invoke}}-1}W^{V}\\ X_{t_{\mathrm{invoke}}:t}(W^{V}\!+\!\Delta^{V})\end{array}\!\!\right]$ (6)
@@
where $X_{1:t_{\mathrm{invoke}}-1}$ and $X_{t_{\mathrm{invoke}}:t}$ are the portions of $X$ coming before and after the aLoRA model is invoked. If $X_{1:t_{\mathrm{invoke}}-1}$ is associated with tokens generated or prefilled by the base model, then $X_{1:t_{\mathrm{invoke}}-1}W^{K}$ and $X_{1:t_{\mathrm{invoke}-1}}W^{V}$ are already in the KV cache and do not need to be recomputed. Similarly, any tokens that have been prefilled or generated by the aLoRA model have keys and values processed by the adapted weights, so $X_{t_{\mathrm{invoke}}:t}(W^{K}+\Delta^{K})$ and $X_{t_{\mathrm{invoke}}:t}(W^{V}+\Delta^{V})$ are also available (except for the current token being generated). As a result, the aLoRA architecture can seamlessly reuse the existing base model KV cache as well as continue to maintain its own KV cache as it generates. Note that adaptations for the MLP blocks that preserve this KV cache property can be done via directly corresponding equations as for the $Q,K,V$ blocks above.
We can formalize this cache reuse claim as follows (proved in the appendix):
Proposition 1 (KV equivalence and aLoRA inference).
For the causal decoder-only transformers we consider, the keys and values (actually, all internal states) prior to $t_{\mathrm{invoke}}$ are identical for the base model and any aLoRA adapter model using (6). Specifically, $K^{base}_{1:t_{\mathrm{invoke}}-1}=K^{adapter}_{1:t_{\mathrm{invoke}}-1}$, and $V^{base}_{1:t_{\mathrm{invoke}}-1}=V^{adapter}_{1:t_{\mathrm{invoke}}-1}$. Inference with the aLoRA adapted model can be done causally one token at a time (by simply increasing $t$ in (6) iteratively) with KV cache reuse.
Furthermore, we can quantify the computation and memory savings:
Proposition 2 (aLoRA vs. LoRA inference costs).
Consider invoking an adapter with $T_{cache}$ tokens of input for which a base model KV cache exists and $T_{new}\ll T_{cache}$ input tokens without cache. [^3] We presume that $T_{new}$ is the invocation sequence and following. The first token generated by the aLoRA adapter requires $O(T_{cache}T_{new})$ operations, while LoRA requires $O((T_{cache}+T_{new})^{2})$. Furthermore, aLoRA must maintain only $O(T_{new})$ additional KV cache memory, while LoRA requires additional $O(T_{cache}+T_{new})$. If $N$ distinct adapters share the first $T_{cache}$ tokens as input, aLoRA costs become $O(NT_{cache}T_{new})$ and $O(NT_{new})$, while LoRA has $O(N(T_{cache}+T_{new})^{2})$ and $O(N(T_{cache}+T_{new}))$.
As the context length and number of concurrent adapters rise, the computation and memory advantages of aLoRA scale linearly. Note that while methods exist for KV cache compression that can reduce overall memory footprint, e.g. Li et al. (2024), and various methods exist increasing the practical scalability of attention, KV cache reuse is still a major factor in practice as context length increases.
Cache reuse patterns: Importantly, while we typically emphasize the aLoRA adapter reusing cache from the base model, the fact that input cache is interchangeable between the base model and every aLoRA adapter presents multiple opportunities for reuse. Consider three regimes:
  1. 1.
    The adapter reuses cache from the base model, e.g. checking a base model generation.
  2. 2.
    The base model reuses cache created when the aLoRA adapter was invoked, e.g. when an adapter checks an input prior to sending to the base model for full generation.
  3. 3.
    Multiple adapters reuses the same cache, e.g. when checking base model outputs on multiple axes. This regime creates the most dramatic advantage for aLoRA over LoRA.
In the results, we explore concrete intrinsics which follow these patterns. Note that any KV cache created by aLoRA adapter for their own generations are not reusable by other adapters or the base model, as these tokens come after the adapter weights are turned on. If these strings needs to be input into other adapters or the base model, they should be prefilled again. This is a fairly minor concern, however, as intrinsics generations are often short and per token, prefill is much faster than generation.
@@
Invocation: We found it useful to demarcate the activation of the aLoRA adapter via a short invocation token sequence. Advantages are elaborated on in Appendix A include allowing more space for the adapted weights to process the input. Typically, this sequence will be appended to the prompt prior to generation, just as a “generation prompt” is typically appended to prompts for instruct-tuned model (e.g. <|assistant|> or some other string). While the aLoRA will often be invoked programmatically, this design in principle allows for the base model (or other intrinsics) to call the aLoRA model themselves (we do not explore this behavior in the current work).
2.1 Training
The activated LoRA framework is explicitly designed for tuning the LLM to modify its output conditioned on an input. This regime matches (but is not limited to) instruction finetuning tasks, where the context tokens are excluded from the loss while finetuning the adapter to produce a specified output sequence (supervised finetuning, or SFT). aLoRA can in principle also be used in RL training pipelines, though we do not explore these in the current work.